• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Jack

    Log entry time 19:03:05 on June 05, 2007

    Entry number 204369

    keyword=Work during the access

    I cycled the power on the camera under the Right Dipole. Last I heard, it was doing ok. I put some kapton tape over the sense resistors on the LQ2 and LQ3 EPCT modules for the current sensing. The thought is it may help to keep the current readback stable. I started gas flowing in the SC3 and SC4 on the Left Detector Stack. SC3(FPP3) had the rotameter bead at 52 and SC4(FPP4) had the bead at 68. Scot and I replaced both of the motor controllers for the Left Shield House doors. We had some confusion in the process, but finally got things working enough to close the doors. I had wanted to look at the threshold power supply for the upper VDC on the Right Arm, but the work on the Shield House doors forced me to postpone that task till later. The voltage readback is drifting around. It has apparently been doing this for a long time.